OpenStudy (anonymous):

how would one solve the equation (dy/dt = -5y +sin(3t)) ?

OpenStudy (bahrom7893):

do u know what to use for them? integrating factors.... etc.. what's the topic?

OpenStudy (bahrom7893):

i dont wanna look for the method...

OpenStudy (bahrom7893):

okay i gotta use integrating factors

OpenStudy (bahrom7893):

dy/dt = -5y + Sin(3t)

OpenStudy (anonymous):

by method of undetermined coefficients if possible

OpenStudy (bahrom7893):

dy = -5y dt + Sin3t dt -5y dt + Sin3t dt - dy = 0 (-5y + Sin3t)dt = dy

OpenStudy (bahrom7893):

(-5y + Sin3t)dt - dy = 0 (5y - Sin3t)dt + dy = 0 Now let M = (5y - Sin3t) and N = 1

OpenStudy (bahrom7893):

My (derivative of M w/h respect to y) is = 5 Nx = 0

OpenStudy (bahrom7893):

(My - Nt)/N = (5-0)/1 = 5 (By the way Nx must be Nt too above)

OpenStudy (bahrom7893):

So let's multiply everything in this d.e. by an integrating factor: myu = e^(Integral of 5dx) = e^(5x)

OpenStudy (bahrom7893):

(5e^(5x)y - e^(5x)Sin3t)dt + e^(5x)dy = 0

OpenStudy (bahrom7893):

Now My = Nt = 5e^(5x)

OpenStudy (bahrom7893):

wow sorry about that, i keep confusing t with x, everywhere x means t, now let me just switch to x, cuz i keep getting mixed up... t is same as x, okay?

OpenStudy (anonymous):

ok

OpenStudy (bahrom7893):

So now Fx ( x ; y ) = M = (5e^(5x)y - e^(5x)Sin3)dx (x is same as t)

OpenStudy (bahrom7893):

sorry forgot Sin(3x)

OpenStudy (bahrom7893):

F = Integral of M with respect to x.

OpenStudy (bahrom7893):

\[F = \int\limits_{}^{}5*e^{5x}y - e^{5x}Sin(3x) dx\]

OpenStudy (bahrom7893):

Sorry i'll have to leave now, so just to finish this, integrate that with respect to x to get something like: e^(5x)y -.................. +h(y)

OpenStudy (bahrom7893):

then you will have: F = e^(5x)y -.................. +h(y). Differentiate the expression you get with respect to y to get: Fy = e^(5x) - or plus ............... +h'(y) Set Fy = N and solve for h'(y) then integrate h'(y) to find h(y) and plug that into F(x;y)

OpenStudy (bahrom7893):

hey u still here? when is this due... i can email it 2 u 2morrow...

OpenStudy (bahrom7893):

HELLO?

OpenStudy (anonymous):

hey sorry, a friend stopped by my room asking for help on a euler's method thing. i have a test over it tomorrow

OpenStudy (bahrom7893):

hey ima send u a couple of examples that use the same principle that were solved out by me, sorry im too tired right now

OpenStudy (bahrom7893):

okay? gimme ur email...

OpenStudy (bahrom7893):

bpw12 gimme an email, i will send u a couple of solved problems that use the same method.. follow them and solve this the same way, sorry too tired now..

OpenStudy (anonymous):

treeclimberbpw@gmail.com thanks btw.

OpenStudy (bahrom7893):

check ur email

OpenStudy (bahrom7893):

this equation is just like one of those; get it in format Mdx + Ndy = 0

OpenStudy (bahrom7893):

and solve like i did...

OpenStudy (bahrom7893):

okay lemme just grab a coffee and shove my head into cold water to wake up lol

OpenStudy (bahrom7893):

okay workin on it

OpenStudy (bahrom7893):

hey emailed you the solution, please click on become a fan if I helped, thanks!